Balance the each of the following half-reactions and write the net-equation               


Al3+ (aq) + 3e- → Al (s)     and   Sn2+ (aq) + 2e- → Sn(s)


1
Expert's answer
2022-06-09T20:21:03-0400

net-equation:

AlCl3 + Sn = SnCl2 + Al

Half-reaction:

Al3+ + 3 e- => Al

Sn – 2 e- => Sn2+


Balaned reaction:

2AlCl3 + 3Sn = 3SnCl2 + 2Al